OpenStudy (anonymous):

the sum of 38 and twice a number is 124. find the number show work

OpenStudy (anonymous):

124-38=86 86/2=43

OpenStudy (akashdeepdeb):

Let that number be x So the equation would be the 38 + 2*thatnumber = 124! So 38+2x=124 Thus 2x=124-38=86 Thus x=86/2 Understood? :)